the height in feet above the ground of an arrow can be modeled by y=-16x^2+62x+4. Can the arrow pass over a tree that is 68 feet tall? why or why not?
Answer by lwsshak3(11628)

the height in feet above the ground of an arrow can be modeled by y=-16x^2+62x+4. Can the arrow pass over a tree that is 68 feet tall? why or why not?
***
y=-16x^2+62x+4
complete the square:
y=-16(x^2-(62/16)x+(62/32)^2)+16(62/32)^2+4
y=-16(x-(62/32))^2+16(62/32)^2+4
y=-16(x-1.9375)^2+64.0625
This is an equation of a parabola that opens downwards with vertex at(1.9375, 64.0625)
maximum height of the arrow≈64 ft
therefore, arrow cannot pass over a tree that is 68 feet tall